I must find the horizontal and vertical asymptotes of : f(x) = x/ x^2 - 9
Found 2 solutions by lwsshak3, lynnlo:
Answer by lwsshak3(11628)

I must find the horizontal and vertical asymptotes of : f(x) = x/ x^2 - 9
..
When the degree of the numerator is 1 less than the degree of the denominator, as in this case, the horizontal asymptote is the x-axis or y=0.
Vertical asymptotes are found by setting the denominator=0, then solving for x-values which makes the function undefined.
x^2-9=0
(x+3)(x-3)=0
x≠3
and
x≠-3
horizontal asymptotes: x-axis
Vertical asymptotes: x=3, x=-3)

Answer by lynnlo(4176)

VERTICAL======================X=0
HORIZONTAL=============Y=-9